If the denominator of 3/5 is increased by a number and the numerator is doubled, the result is 1. Find the number.
Found 2 solutions by addingup, josgarithmetic:
Answer by addingup(3677)

2(3)/x+5 = 1
6/x+5 = 1
take the reciprocal on both sides:
x+5/6 = 1
x+5 = 6
x = 1
~~~~~~~~~~~~~~~~~
Check:
6/1+5 = 1 correct

Answer by josgarithmetic(39630)